OpenStudy (help!!!!):

23. What is the fractional equivalent of 12.5%? A. 1/4 B. 2/9 C. 1/5 D. 1/8 E. 2/7 How did you find the answer?

OpenStudy (anonymous):

12 = 12/1 and .5 = 1/2, they are added together in the form 12.5, the % sign denotes that the number is over 100 so you would have (12/1 + 1/2)/100

OpenStudy (help!!!!):

Why divide by a hundred

OpenStudy (anonymous):

because of the % sign

OpenStudy (anonymous):

12.5% = 12.5/100

OpenStudy (mathstudent55):

A percent is always a ratio based on 100. 12.5% means 12.5 out of 100 or 12.5/100 Now multiply the numerator and denominator by 2 to get whole numbers, 25/200, and reduce the fraction. 25 and 200 are both divisible by 25.